OpenStudy (anonymous):

A rock is dropped from a sea cliff and the sound of it striking the ocean is heard 5.5 s later. If the speed of sound is 340 m/s, how high is the cliff?

OpenStudy (anonymous):

1/2gt^2 is the distance we are trying to find. Isolating the time, we get:\[\sqrt{\frac{2h}{g}}=t\]The total time is:\[5.5=t+t_s=\sqrt{\frac{2h}{g}}+\frac{h}{v_s}\]Isolate h and you get your answer.